A Food Marketing Institute found that 42% of households spend more than $125 a week on groceries. Assume the population proportion is 0.42 and a simple random sample of 55 households is selected from the population. What is the probability that the sample proportion of households spending more than $125 a week is between 0.38 and 0.5?

Note: You should carefully round any z-values you calculate to 4 decimal places to match wamap's approach and calculations.

Answer = (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Solution:

mean=p=0.42

standard deviation=sqrt(p*(1-p)/n=sqrt(0.42*(1-0.42)/55)=0.06655142

P(0.38<p^<0.5)

z=p^-p/sqrt(p*(1-p)/n)

P(0.38-.42/0.06655142<Z<0.5-0.42/0.06655142)

P(-0.601039<Z<1.202078)

P(-0.6010<Z<1.2021)

P(Z<1.2021)-P(Z<-0.6010)

=0.8853-0.2739

=0.6114

ANSWER:

0.6114